# On the Effectiveness and Generalization of Race Representations for Debiasing High-Stakes Decisions

### v1

@@ 1 Introduction
To that end, we first introduce synthetic tasks and datasets, Admissions and Hiring, in which models are given hypothetical applicant profiles and are asked whether to accept or reject them. We assess models’ biases by giving them applicant names that are highly suggestive of their race, and measure the disparity in outcomes across races using our novel fairness metric, BiasScore. Working with Gemma 2B Instruct and LLaMA 3.2 3B Instruct, we find that models indeed exhibit racial biases in Admissions and Hiring. In both decision tasks, Gemma favors White and Asian applicants over Black and Latino applicants, and LLaMA favors Asian and Latino applicants over Black and White ones.
@@
Figure 1: We consider two decision tasks: Admissions and Hiring, and examine two approaches to control model behavior. In prompt engineering, we attempt to debias by adding various instructions. To modify model internals, (1) we learn a “race subspace” to intervene so that the original decision becomes the counterfactual decision (i.e., as if the race inferred from the name is the same as that from the source run). (2) We debias model decisions by averaging the representation in the race subspace across a batch of samples, which removes the variance in race representations between applicants.
@@
Our first approach to debias model decisions build on the model’s ability to follow instructions, which has proven effective in many cases (Brown et al., 2020; Wei et al., 2022; Tamkin et al., 2023; Tseng et al., 2024). However, we find that different prompting strategies to promote fairness end up exacerbating the biases or even completely derailing model generation. For example, asking Gemma to not discriminate based on sensitive attributes in Hiring can increase its BiasScore by 137%, and doing so for LLaMA can cause it to indiscriminately accept all applicants. Thus, prompt engineering can unpredictably alter model behavior in high-stakes decision-making, rendering it unreliable.
@@
Seeking a more principled and effective approach, we attempt to pinpoint the mechanism causing biased decisions in LLMs and modify it to mitigate biases. We use distributed alignment search (Geiger et al., 2024; Wu et al., 2024b) to identify a subspace in model hidden representation dedicated to encoding race, a “race subspace”, with 74-85% interchange intervention accuracy (see exact definitions in Geiger et al. (2024) and Section 5) indicating strong race representation. Hence, models implicitly keep track of the applicant’s race despite it not being given in their profile.
@@
We debias model decisions by averaging the values of all races in the subspace, thereby erasing the variance between different races, or by projecting out the race subspace from the model’s hidden representation, which “deletes” race information altogether. Compared to prompt engineering, representation-based debiasing are much more effective. Race averaging reduces Gemma’s bias by 37-57% and LLaMA’s bias by 32% in Admissions. This suggests that mechanistic approaches are promising for mitigating biases in LLM decisions.
@@
We will release code and data upon publication.
@@ 3 Decision tasks and experiment setup
[…]
@@ 4 Prompting is Ineffective at Debiasing
We show Gemma and LLaMA can be racially biased in their admissions and hiring decisions. Given their instruction-following ability, we attempt to mitigate these biases through prompting. However, this fails to promote fairness even across a wide range of strategies.
@@
Debiasing through prompting. Following Tamkin et al. (2023), we explicitly prompt models for fairness using various strategies. We use the same prompt template as above and append a fairness prompt sentence at the end according to the following strategies: Simple: “Do not discriminate against applicants based on race, gender, or religion.” No Affirmative Action: “We do not believe in nor support affirmative action and so it should have no bearing on your decision.” “Very” (×$\times k$) important: “It is very important that race, gender, religion, and other demographic characteristics do not influence this decision.” The word “very” is repeated $k$ times for $k=1,2,4$. Illegal: “Do not discriminate against applicants based on race, gender, or religion as it is illegal and unethical.”
@@ 5 Learned representations enable effective debiasing
In this section, we use causal abstraction (Geiger et al., 2023) to identify representations of race in models’ hidden activations, on which we can intervene to mitigate the identified racial biases.
@@ 5.1 Finding the Race Subspace
The key intuition of causal abstraction is to search for a mapping between neural representations and a causal model. If successful, interventions on the neural representations achieve similar effects on the outcome variable of interest as interventions on the causal model. Here we provide an overview and connect it with our context (see a high-level illustration in Figure 1). For full details on the theory of causal abstraction, please refer to Geiger et al. (2023; 2024); Wu et al. (2024b).
Causal model. Formally, let $\mathcal{C}$ be a causal model defined as a set of input variables $\mathbf{U}$, intermediate variables $\mathbf{V}$, and functions $\{f_{V}\mid V\in\mathbf{V}\}$ that assign values to every intermediate variable $V$ using the values of its parents. For example, $\mathcal{C}=(A=3)\land(B=5)$ is a causal model with input variables $A$, $B$ and intermediate variables $(A=3)$, $(B=5)$, and $\mathcal{C}$. In our case, a relevant causal model is one that describes the model’s decision-making process, such as $\mathcal{C}=(\text{Race}=\dots)\land(\text{GPA}=\dots)\land\dots$ in Admissions.
Often, there can be multiple causal models that fully explain a model’s decision or an algorithm, such as multiple ways to sort a list of numbers. Since we only care about the race representation and do not want to construct arbitrary decision boundaries on the qualification variables, we use the LLM itself as the causal model, denoted as $\mathcal{N}$.
Interchange intervention. Let $\{(s_{i},t_{i})\}_{i=1}^{n}$ be source and target input pairs to the causal model $\mathcal{N}$, where each $s_{i}$ and $t_{i}$ is a tuple of input variables defined in Table 1. $s_{i}=(Q^{s}_{i},N^{s}_{i})$, where $Q^{s}_{i}$ is the set of qualifications and $N^{s}_{i}$ is the name in $s_{i}$. Similarly, $t_{i}=(Q^{t}_{i},N^{t}_{i})$. Note that $Q^{s}_{i}$ can be different from $Q^{t}_{i}$. Let $V$ be the intermediate variable that we want to find a neural representation for: an applicant’s race inferred from their name. An interchange intervention $\textsc{Int}(\mathcal{N},V,s_{i})$ returns a modified $\mathcal{N}$, where the race is changed from that in $t_{i}$ to that in $s_{i}$. $\textsc{Int}(\mathcal{N},V,s_{i})(t_{i})$ is this new model’s output for $t_{i}$. In practice, $\textsc{Int}(\mathcal{N},V,s_{i})(t_{i})$ is computed using $\mathcal{N}(Q^{t}_{i},N^{s}_{i})$, the model’s output given a counterfactual profile with the qualifications from $t_{i}$ and the name from $s_{i}$. See Appendix C for more examples of interchange interventions.
Distributed interchange intervention. We aim to find a subspace in the model’s hidden representation that achieves the same effect as changing the causal variable race. Let $H(\cdot)\in\mathbb{R}^{d}$ be the hidden representation at some target token and layer of the language model $\mathcal{N}$. By abuse of notation, we also use $\{(s_{i},t_{i})\}_{i=1}^{n}$ to represent source and target prompt pairs to the network. However, instead of tuples of variables, $s_{i}$ and $t_{i}$ are now tokens in the prompt. Let $R$ be the subspace corresponding to race. A distributed interchange intervention $\textsc{DII}(\mathcal{N},R,s_{i})(t_{i})$ replaces $H(t_{i})$ with
@@
Distributed alignment search (DAS). Recall that $s_{i}$ and $t_{i}$ refer to both inputs to the causal model and the language model, and the output of both models takes on binary values “Yes” or “No”. Given a model $\mathcal{N}$ and race variable $V$, let $\mathbb{P}(o\mid t_{i}):=\textsc{Int}(\mathcal{N},V,s_{i})(t_{i})$ be the (one-hot) distribution on the outputs of the intervened causal model given $t_{i}$ as input. For a hypothesized race subspace $S$, let $\mathbb{Q}_{S}(o\mid t_{i}):=\textsc{DII}(\mathcal{N},S,s_{i})(t_{i})$ be the distribution on the outputs of the intervened language model given $t_{i}$ as input. Then, we search over subspaces $S$ by minimizing the following objective:
@@
Intuitively, this objective encourages searching for a subspace that achieves the outcome by modifying the activations as if we replaced the input name tokens. We can either search over subspaces of all dimensions or over those of a specified dimension. In our case, we found that subspaces of 500 and 1000 dimensions give the best debiasing performance for Gemma and LLaMA 3.2, respectively.
@@
$\text{IIA}=\frac{\sum_{i=1}^{N}\mathbbm{1}\big{[}\textsc{DII}(\mathcal{N},R,s_{i})(t_{i})=\textsc{Int}(\mathcal{N},V,s_{i})(t_{i})\big{]}}{N},$
where $N$ is the total number of input pairs. The IIA measures how often interventions on the race subspace produce the same output as changing the input name, i.e., the degree to which the subspace $R$ represents the causal variable race.
@@ 5.2 Alignment training results
(a) Gemma’s alignment training results
(b) Gemma’s best IIAs
@@
Figure 3 shows alignment training results for Gemma on the test set (see LLaMa results in Appendix C). At many locations, the IIA is random, suggesting that race is localized to specific tokens and layers. On the last token of the name (token 18), we observe higher IIAs of up to 66%, although they are lower than the IIAs at the final token, which reach up to 85% at layer 10 for Admissions and 84% at layer 12 for Hiring, suggesting that the model derives race information from applicants’ names. Similarly, we observe the best IIA of up to 72% for LLaMA 3.2 at the final prompt token, layers 24-25 (Appendix C). This is likely due to padding within a batch (as names and universities have different lengths) causing the name tokens to misalign between the source and target, i.e., some token positions that represent race in the source may not do so in the target.
In contrast, the final token representation will always include race (if it is a factor of consideration) because it directly affects the next predicted token via a linear transformation and softmax. Note that, due to this property, the last-token representation has garnered special attention in the interpretability literature. Li et al. (2024); Wang & Veitch (2024) intervene on attention head ouputs at this location to increase models’ truthfulness, and Park et al. (2023) investigates the geometry of its residual stream. Therefore, when we perform debiasing in Section 5.3, we work with the last-token representations at layers with the best IIAs. For more details on alignment training results, please refer to Appendix C.
@@ 5.3 Representation-based debiasing results
Debiasing through interventions. We use the race subspace identified in Section 5.2 and experiment with the following interventions. Race Averaging: we replace the race representation with the average representation over all races. Race Projection: we project out the race subspace, effectively removing race from the model’s consideration. Accordingly, our baselines include: Full Averaging: we replace the entire model’s representation at a layer with the average representation over a batch. Random Projection: we project out a random 500-dimensional subspace.
Interventions outperform prompting in debiasing Gemma. In Table 3, Gemma’s original BiasScore in Admissions is 13.54. The model prefers Asian and White applicants and discriminates against Black and Latino applicants. Both Race Averaging and Projection effectively reduce the bias. Race Averaging reduces the BiasScore by 57.53% from 13.54 to 5.75 while only increasing the average acceptance rate by 5% absolute. Race Projection more effectively reduces the bias, by 80.8%, but it increases the average acceptance rate slightly more than Race Averaging, by 7.56% absolute. These changes in the outcome are acceptable because the acceptance rates of the Black and Latino applicants must increase to match those of White and Asian applicants. In contrast, the baselines fail to reduce the bias, where Full Average seems to leave Gemma’s decisions unchanged and Random Projection results in all rejections, a behavior we also observe in many fairness prompts.
The results are similar in Hiring: Race Averaging effectively reduces the bias by 37.89% while only increasing the average acceptance rate by 2.31% absolute, which reflects the increase in minority acceptances. It is much more effective than averaging the whole layer 12 representation, which increases the BiasScore by 242%. However, Race Projection fails to work as it results in all rejections, similar to a random projection. Therefore, Race Averaging is the overall most effective method for debiasing Gemma.
We successfully reduced LLaMA’s bias in Admissions but not Hiring. Please refer to Appendix D for more details.
6 Do Race Representations Generalize?
There is evidence in mechanistic interpretability for certain concepts being generalized across different contexts. Arditi et al. (2024) ablate a single “refusal direction” across all layers and token positions to bypass models’ safety mechanism. Anthropic (2024) activates a set of features related to the Golden Gate Bridge to get Claude to mention it in all conversations. Given this evidence and our promising debiasing results, we examine whether the race subspace in Admissions is also universal and generalizes across different settings. To our surpirse, we find this to not be the case.
@@ 6 Do Race Representations Generalize?
Generalization experiment setup. Our goal is to test if an optimal race subspace trained on Admissions can be used to debias other decisions settings. We consider three different settings: cross-prompt, cross-task family, and cross-explicitness. For cross-prompt, we format an applicant’s profile in a bulleted list instead of free text. We refer to this analysis as free-text->list. For cross-task family, we investigate generalization between Admissions and Hiring (Admissions->Hiring). For cross-explicitness, we consider a different setting of Admissions where race is explicitly provided in the profile (implicit->explicit). Examples of prompts with bulleted list format and explicit race mentions can be seen in Appendix A.
@@
These limited results reflect the overall low cross-setting IIA (Figure 15, Appendix E). Admissions->Hiring fails to outperform the optimal subspace trained on Hiring, free-text->list fails to outperform a random subspace, and implicit->explicit fails to outperform random. We note that achieving effective debiasing is generally easier than achieving high IIA. In debiasing, we only need to perturb a part of the race subspace to affect the race representation. In contrast, a valid interchange intervention requires a higher degree of exactness for the source representation to stay within the model’s distribution.
We observe even more limited generalization in LLaMA 3.2, where both Race Averaging and Race Projection fail to meaningfully debias the target settings. We attribute this worse performance to LLaMA’s weaker race representation, where the best IIA achieved is only 72%. Our results suggest that race representations are specific to the decision task setting. Future work on debiasing will likely have to tailor strategies to each studied setting.

### v2

@@ 1 Introduction
To that end, we first introduce synthetic tasks and datasets, Admissions and Hiring, in which models are given hypothetical applicant profiles and are asked whether to accept or reject them. We assess models’ biases by giving them applicant names that are highly suggestive of their race, and measure the disparity in outcomes across races using our novel fairness metric, BiasScore. Working with Gemma 2B Instruct and LLaMA 3.2 3B Instruct, we find that models exhibit racial biases in Admissions and Hiring. In both decision tasks, Gemma favors White and Asian applicants over Black and Latino applicants, and LLaMA favors Asian and Latino applicants over Black and White ones.
@@
Figure 1: We consider two decision tasks: Admissions and Hiring, and examine two approaches to control model behavior. In prompt engineering, we attempt to debias by adding various instructions. To modify model internals, (1) we learn a “race subspace” to intervene so that the target decision becomes the counterfactual decision (i.e., as if the inferred race is the same as that from the source text). (2) We debias model decisions by averaging the representation in the race subspace across a batch of samples, which removes the variance in race representations between applicants.
@@
Our first approach to debiasing model decisions builds on the model’s ability to follow instructions, which has proven effective in many cases (Brown et al., 2020; Wei et al., 2022; Tamkin et al., 2023; Tseng et al., 2024). However, we find that different prompting strategies to promote fairness end up exacerbating the biases or even completely derailing model generation. For example, asking Gemma to not discriminate based on sensitive attributes in Hiring can increase its BiasScore by 137%, and doing so for LLaMA can cause it to indiscriminately accept all applicants. Thus, prompt engineering can unpredictably alter model behavior in high-stakes decision-making, rendering it unreliable.
@@
Seeking a more principled and effective approach, we attempt to pinpoint the mechanism causing biased decisions in LLMs and modify it to mitigate biases. We use distributed alignment search (Geiger et al., 2024; Wu et al., 2024b) to identify a subspace in models’ hidden representations dedicated to encoding race, a “race subspace”, with 74-85% interchange intervention accuracy (see exact definitions in Geiger et al. (2024) and Section 5) indicating strong race representation. Hence, models implicitly keep track of the applicant’s race despite it not being given in their profile.
@@
We debias model decisions by averaging the values of all races in the subspace, thereby erasing the variance between different races, or by projecting out the race subspace from the model’s hidden representation, which “deletes” race information altogether. Compared to prompt engineering, representation-based debiasing is much more effective. Race averaging reduces Gemma’s bias by 37-57% and LLaMA’s bias by 32% in Admissions. This suggests that mechanistic approaches are promising for mitigating biases in LLM decisions.
@@
We release code and data at https://github.com/ChicagoHAI/llm-prediction-bias.
@@ 3 Decision Tasks and Experiment Setup
Metrics for fairness. Although the notion of fairness can depend on the context, in this work, we treat fairness as equality of opportunities, where, given the same qualifications, we expect the same decision for all races. We consider two fairness metrics. BiasScore measures the equality in acceptance rates across races: $100\times\frac{\sum_{i=1}^{N}\sigma(\{D(q_{i},r)\ \mid\ r\in\operatorname{Race}\})}{N}$, where $\sigma$ denotes the standard deviation, $D\in\{0,1\}$ is the model’s decision for the profile of qualifications $q_{i}$ and race $r$, NN is the number of profiles, and $\operatorname{Race}=\{\text{Asian, Black, Latino, White}\}$. Since the decision is either 0 or 1 for each applicant, the BiasScore ranges from 0 (when there is no inequality) to 50 (when there are two races favored and two disfavored).
While reducing BiasScore is the main goal, methods must also minimally change the average acceptance rate. We measure this with Outcome $\Delta$, the change in the average acceptance rate after an intervention. In addition, we report $\mathbb{P}(\text{``Yes''}\mid\text{race}$) for each method to observe the effect on overall acceptance rates.
@@ 4 Prompting is Ineffective at Debiasing
In this section, we show Gemma and LLaMA can be racially biased in their admissions and hiring decisions. Given their instruction-following ability, we attempt to mitigate these biases through prompting. However, a wide range of strategies all fail to promote fairness.
@@
Debiasing through prompting. Following Tamkin et al. (2023), we explicitly prompt models for fairness using various strategies. We use the same prompt template as above and append a fairness prompt sentence at the end according to the following strategies: Simple: “Do not discriminate against applicants based on race, gender, or religion.” No Affirmative Action: “We do not believe in nor support affirmative action and so it should have no bearing on your decision.” “Very” ($\times k$) important: “It is very important that race, gender, religion, and other demographic characteristics do not influence this decision.” The word “very” is repeated $k$ times for $k=1,2,4$. Illegal: “Do not discriminate against applicants based on race, gender, or religion as it is illegal and unethical.”
Due to a lack of groundtruth label, we cannot perform a fair comparison of the above approaches with few-shot and chain-of-thought prompting. We choose to omit them from Table 2 and refer the reader to Appendix B for those results. In an nutshell, few-shot and chain-of-thought prompting similarly fail to mitigate biases, resulting in a large Outcome $\Delta$.
@@ 5 Learned Representations Enable Effective Debiasing
To pursue an alternative mechanistic approach, we use causal abstraction (Geiger et al., 2023) to identify representations of race in models’ hidden activations, on which we can intervene to mitigate the identified racial biases.
@@ 5.1 Finding the Race Subspace
The key intuition of causal abstraction is to search for a mapping between neural representations and a causal graph. If successful, then interventions on the neural representations achieve similar effects on the outcome variable of interest as interventions on the causal graph. Here we provide an overview and connect it with our context (see a high-level illustration in Figure 1). For full details on the theory of causal abstraction, please refer to Geiger et al. (2023; 2024); Wu et al. (2024b).
Formally, let $\mathcal{C}$ be a causal model as defined by Pearl et al. (2016) and $V$ be the variable we want to intervene on (race). Let $\{(s_{i},t_{i})\}_{i=1}^{n}$ be source and target input pairs to the causal model. An interchange intervention $\textsc{Int}(\mathcal{C},V,s_{i})$ returns a modified $\mathcal{C}$, where the race is set to that induced by the name in $s_{i}$. $\textsc{Int}(\mathcal{C},V,s_{i})(t_{i})$ is this new model’s output for $t_{i}$. Since $\textsc{Int}(\mathcal{C},V,s_{i})(t_{i})$ and $\mathcal{C}(t_{i})$ are minimally different in the race, any change in the output can be attributed to the change in race. The causal model in our work is the LLM itself. We aim to find a subspace that achieves the outcome as if we replaced the input name.
A distributed interchange intervention is the neural counterpart of the interchange intervention, which acts on subspaces in neural representations instead of discrete variables. Let $\mathcal{N}$ be a neural network and $H(v)\in\mathbb{R}^{d}$ be the hidden representation at some target token and layer. Similarly, $\{(s_{i},t_{i})\}_{i=1}^{n}$ are source and target prompt pairs to the network. Most importantly, let $R$ be the subspace corresponding to our desired variable, race. A distributed interchange intervention $\textsc{DII}(\mathcal{N},R,s_{i})(t_{i})$ replaces $H(t_{i})$ with
@@
Distributed alignment search (DAS). We abuse notation and let $s_{i}$ and $t_{i}$ refer to both inputs to the causal model and the language model. Recall that the output variable is $o$ that takes on binary values “Yes” or “No”. Given a causal model $\mathcal{C}$ and race variable $V$, let $\mathbb{P}(o\mid t_{i}):=\textsc{Int}(\mathcal{C},V,s_{i})(t_{i})$ be the (one-hot) distribution over the outputs of the intervened causal model given $t_{i}$ as input. This serves as the groundtruth signal. For a language model $\mathcal{N}$ and a hypothesized race subspace $S$, let $\mathbb{Q}_{S}(o\mid t_{i}):=\textsc{DII}(\mathcal{N},S,s_{i})(t_{i})$ be the distribution over the outputs of the intervened language model given $t_{i}$ as input. Then, we search over subspaces $S$ by minimizing the following objective:
@@
We can either search over subspaces of all dimensions or over those of a specified dimension. In our case, we found that subspaces of 500 and 1000 dimensions give the best debiasing performance for Gemma and LLaMA 3.2, respectively.
@@
$\text{IIA}=\frac{\sum_{i}\mathbbm{1}\big[\textsc{DII}(\mathcal{N},R,s_{i})(t_{i})=\textsc{Int}(\mathcal{C},R,s_{i})(t_{i})\big]}{N},$
where $N$ is the total number of input pairs. The IIA measures how often interventions on the language model produce the same output as corresponding interventions on the causal model, i.e., the degree to which the subspace $R$ represents the causal variable race.
@@ 5.2 Alignment Training Results
(a) Gemma’s alignment training results.
(b) Gemma’s best IIAs.
@@
Figure 3 shows alignment training results for Gemma on the test set (see LLaMA results in Appendix C). At many locations, the IIA is random, suggesting that race is localized to specific tokens and layers. On the last token of the name (token 18), we observe higher IIAs of up to 66%, although they are lower than the IIAs at the final token, which reach up to 85% at layer 10 for Admissions and 84% at layer 12 for Hiring, suggesting that the model derives race information from applicants’ names. Similarly, we observe the best IIA of up to 72% for LLaMA 3.2 at the final prompt token, layers 24-25 (Appendix C). The lower IIAs on name tokens are likely due to padding within a batch (as names and universities have different lengths) causing the name tokens to misalign between the source and target, i.e., some token positions that represent race in the source may not do so in the target.
In contrast, the final token representation will always include race (if it is considered) because it directly affects the next predicted token via a linear transformation and softmax. Note that, due to this property, the last-token representation has garnered special attention in the interpretability literature. Li et al. (2024); Wang & Veitch (2024) intervene on attention head outputs at this location to increase models’ truthfulness, and Park et al. (2023) investigates the geometry of its residual stream. Therefore, when we perform debiasing in Section 5.3, we work with the last-token representations at layers with the best IIAs. For more details on alignment training results, please refer to Appendix C.
@@ 5.3 Representation-based Debiasing Results
Debiasing through interventions. We use the race subspace identified in Section 5.2 and experiment with the following interventions. Race Averaging: we replace the race representation with the average representation over all races. Race Projection: we project out the race subspace, effectively removing race from the model’s consideration. Our baselines are: Full Averaging: we replace the entire model’s representation at a layer with the average representation over a batch. Random Projection: we project out a random 500-dimensional subspace. Given the lack of a true counterfactual scenario where the model is entirely fair, we lack a principled way to interpret the Outcome $\Delta$. We approximate the model’s “neutral” behavior by omitting the applicant’s name from the prompt, indicated by No Name in Table 3.
Interventions outperform prompting in debiasing Gemma. In Table 3, Gemma’s original BiasScore in Admissions is 13.54. The model prefers Asian and White applicants and discriminates against Black and Latino applicants. Both Race Averaging and Projection effectively reduce the bias. Race Averaging reduces the BiasScore by 57.53% from 13.54 to 5.75 while only increasing the average acceptance rate by 5% absolute. Race Projection more effectively reduces the bias, by 80.8%, but it increases the average acceptance rate slightly more than Race Averaging, by 7.56% absolute. In both cases, Outcome $\Delta$ is lower than that in No Name. In contrast, the baselines fail to reduce the bias: Full Averaging leaves Gemma’s decisions unchanged and Random Projection results in all rejections, a behavior we also observe in many fairness prompts.
The results are similar in Hiring: Race Averaging effectively reduces the bias by 37.89% while only increasing the average acceptance rate by 2.31% absolute, which reflects the increase in minority acceptances. It is much more effective than averaging the whole layer 12 representation, which increases the BiasScore by 242%. However, Race Projection fails to work as it results in all rejections, similar to a random projection. Therefore, Race Averaging is the overall most effective method for debiasing Gemma. We successfully reduced LLaMA’s bias in Admissions but not Hiring. Please refer to Appendix D for more details.
6 Race Representations Fail to Generalize across Prompts
There is evidence in mechanistic interpretability for certain concepts being generalized across different contexts. Arditi et al. (2024) ablate a single “refusal direction” across all layers and token positions to bypass models’ safety mechanism. Anthropic (2024) activate a set of features related to the Golden Gate Bridge to get Claude to constantly mention it in conversations. Given this evidence and our positive debiasing results, we examine whether the race subspace in Admissions is also universal and generalizes across different settings. To our surprise, we find this to not be the case.
@@ 6 Race Representations Fail to Generalize across Prompts
Generalization experiment setup. Our goal is to test if a race subspace trained on Admissions can be used to debias other decision settings. We consider three different settings: cross-prompt, cross-task family, and cross-explicitness. For cross-prompt, we format an applicant’s profile in a bulleted list instead of free text. We refer to this analysis as free-text->list. For cross-task family, we investigate generalization between Admissions and Hiring (Admissions->Hiring). For cross-explicitness, we consider a different setting of Admissions where race is explicitly provided in the profile (implicit->explicit). Examples of prompts with bulleted list format and explicit race mentions can be seen in Appendix A.
@@
These limited results reflect the overall low cross-setting IIA (Figure 15, Appendix E). Admissions->Hiring fails to outperform the optimal subspace trained on Hiring, free-text->list fails to outperform a random subspace, and implicit->explicit fails to outperform random. We note that achieving effective debiasing is generally easier than achieving high IIA. In debiasing, we only need to perturb a part of the race subspace to degrade the race representation. In contrast, a valid interchange intervention requires a higher degree of exactness for the source representation to stay within the model’s distribution.
We observe more limited generalization in LLaMA 3.2, where both Race Averaging and Race Projection fail to meaningfully debias the target settings. We attribute this worse performance to LLaMA’s weaker race representation, where the best IIA achieved is only 72%. Our results suggest that race representations are specific to the decision task setting. Future work on debiasing will likely have to tailor strategies to each studied setting.
Free text List Random Baseline
Free text \ 63.99 64.01
List 63.99 \ 63.96
Random baseline 64.01 63.96 \
Table 5: Similarity scores (closer to 0 is better) between race subspaces trained on different prompt formats and a random baseline. DAS projections use rotation and masking, so the score is the Frobenius norm of the difference between two rotation matrices. The subspace dimension is 500.
Mechanistic evidence for race representations’ prompt-dependence.
Our results suggest that Gemma uses a different race representation for each prompt format or task variation. We test this hypothesis by comparing the similarity between two race subspaces trained on different prompts, free-text and list. The two race subspaces achieve IIAs of 85% and 73% at layer 10, respectively. In practice, the projection to the race subspace is implemented by a rotation followed by neuron masking (Wu et al., 2024b), where column $i$ in the rotation matrix shows the transformation of the $i^{th}$ standard basis vector. We can compare two rotation matrices via the Frobenius norm between their difference, $\lVert A-B\rVert_{F}$, which gives us a way to measure the distance (similarity score) between two subspaces. Table 5 suggests that different prompt formats induce different race subspaces in the model’s representations, as the free-text and list subspaces are no more similar to each other than they are to a random subspace. This raises an interesting question about why models are seemingly “redundant” in their representation of race (and likely other concepts as well), which we leave to future work.
Each prompt induces a unique subspace that generalizes across races and names.
While cross-prompt generalization is limited, within the same prompt, we found the discovered race subspace is unique and canonical–generalizing across unseen races and names. To show name and race-generalization, we created a training set of 80% of the available names, and a test set of the other 20%. The IIA on this test set is 81.39%, indicating name-generalization. For race-generalization, we made a training set of non-Asian names, and a test set of only Asian names. After finding a race subspace using the training set, the IIA on the test set using this subspace is 88%. This suggests that the subspaces may generalize to notions of race and ethnicities beyond the four coarse-grained ones considered in this paper.
We investigate whether DAS can still find a race subspace after we remove one from an earlier layer. In this experiment, we perform debiasing interventions, Race Averaging and Race Projection, at layer 10 in the Admissions task, free-text format. Subsequently, we run DAS on layer 11 onward to see if there exist other causal race subspaces. The results show that after removing the race representation at layer 10, the IIAs at subsequent layers are at or below random (Table 16), suggesting that we have completely removed information about the applicant’s race. So although a model might use different race representations for different prompts, within the same prompt, the representation is canonical and unique.